OpenStudy (aaronandyson):

A bus starts from rest with a constant acceleration of 5ms^-2.At the same time a car travelling at a constant velocity of 50ms^-1 overtakes and passes the bus..How fast was the bus travelling when the two were side by side?

OpenStudy (aaronandyson):

i think the car is behind the bus and the times is 0 sec

OpenStudy (aaronandyson):

Fur bus: v = 5t s = 2.5t^2 for car: s = vt = 50t 2.5t^2 = 50t t = 20 v= 5t=5*20 = 100ms^-2

OpenStudy (radar):

You are probably going to have to convert this data for the bus & car to distance and when the distances are equal they will be side by side.

OpenStudy (radar):

Oh good, that looks like what you did!

OpenStudy (radar):

I don't think the velocity is expressed in the correct unit just m/s not m/s^2.

OpenStudy (aaronandyson):

oops,typo. is my method right/correct?

OpenStudy (mayankdevnani):

when t = 20 s, both the car and the bus went : s = 50(20) m = 1000 m or 1 km at that time, the bus went at 5t or 100 m/s . Yep! you're correct

OpenStudy (radar):

@AaronAndyson good work.